OpenStudy (anonymous):

Two different chemotherapy regimens can be used for the treatment of breast cancer. The treatment with doxorubicin is known to show a remission rate 40% of the time. This treatment costs $1,000. The other treatment, with docetaxel, shows a remission rate 25% of the time and costs $850. The two treatment plans are as follows: Plan A: Treatment with doxorubicin—if not effective, treatment with docetaxel. Plan B: Treatment with docetaxel—if not effective, treatment with doxorubicin. If you are the patient, which statement is the best choice? PLEASE HELP

OpenStudy (amistre64):

i never was good at these subjective problems ... i recall that expected value is mean times cost if you want to weigh the respective dollar amounts for paying for an ineffective treatment

OpenStudy (anonymous):

is there any type of formula like standard deviation or somethin like that

OpenStudy (amistre64):

.25(850) = 212.50 .40(1000) = 400.00 but im not sure if the standard error we did on the last one would be useful or not .... since we dont have a sample size

OpenStudy (anonymous):

A. Based on the probability of overall remission rate, plan A should be selected over plan B. Based on the probability of overall remission rate, plan B should be selected over plan C. Based on the cost of the first treatment alone, plan A should be selected over plan B. D. Based on the overall cost of treatment, plan B should be selected over plan A. E. Based on the probability of survival rate and the cost of treatment, both plans are equivalent, so either can be selected.

OpenStudy (anonymous):

The answers dont even have numbers in them...

OpenStudy (anonymous):

does that mean that the one that cost 1000$ would be less effective since it cost more money

OpenStudy (amistre64):

to me, E is not accurate since they have different expectation values

OpenStudy (anonymous):

docetaxel is plan B and it cost less so wouldnt it be D

OpenStudy (amistre64):

as a patient, you would prolly care more about success rate then cost.

OpenStudy (amistre64):

this reminds me of those type 1 or 2 error things

OpenStudy (anonymous):

i never realy thought about that, thats true

OpenStudy (amistre64):

my opinion is: as a patient with cancer, you would want the one with the best success rate. But i cant guarentee that as the correct option on a test or quiz :)

OpenStudy (anonymous):

its alright, i'd rather have your opinion than mine, ive failed this semester exam like 5 times already lol

OpenStudy (amistre64):

as a money pinching old man ... try the cheap one first, if you win yay, if not ... dole out for the other one ;)

OpenStudy (anonymous):

alright thank you (:

OpenStudy (amistre64):

good luck

OpenStudy (chaosguardian14):

to calculate the expected cure rate for both plans use the equation: prob of cure for plan a= p(a cured it)+(b cured it)*(a didnt cure it) prob of cure for plan b= p(b cured it)+(a cured it)*(b didnt cure it) to calculate the expected cost of the plans use expected cost of a= cost(a)+cost(b)*(a didnt cure it) expected cost of b= cost(b)+cost(a)*(b didnt cure it) i know this question was asked 2 years ago but this is for those who r like me and just google questions and just so happen to find this one